# arxiv-appendix

### v1

@@ 3.3 Methodology
$r_{k,s}^{t,j}:=\sqrt{\frac{C_{n}\sum_{m=0}^{M}(\lambda_{t,k,s}^{m,j})^{2}\left\{\left(\varepsilon_{m}^{-2}t_{s}^{m,j}\right)\vee\sum_{i=1}^{T_{m}(t)}\tilde{U}_{i,k,s}^{\mathrm{Q}_{m},j}a_{i,s}^{m,j}\right\}}{\left(\sum_{m=0}^{M}\lambda_{t,k,s}^{m,j}\sum_{i=1}^{T_{m}(t)}\tilde{U}_{i,k,s}^{\mathrm{Q}_{m},j}a_{i,s}^{m,j}\right)^{2}}}$ (28)
@@
where $C_{n}\asymp\log(n)$ is specified in the proof. As shown in LABEL:lem:finitedifferenceboundauxilary, (28) provides a confidence bound applicable to the multiple-source estimator, with a rationale similar to that of (14). Note that the term $\sum_{i=1}^{T_{m}(t)}\tilde{U}_{i,k,s}^{\mathrm{Q}_{m},j}a_{i,s}^{m,j}$ approximately corresponds to the number of samples falling in $B_{s}^{j}$ while pulling arm $k$. This quantity generally increases with $\kappa_{m}$ and decreases with $\gamma_{m}$, in view of the definitions of these quantities. Therefore, as a statistic, (28) encodes information about $\kappa_{m}$ and $\gamma_{m}$, which is a key factor enabling our estimator to adapt to these parameters.
Given the estimator (26) and the confidence bound (28), we define the bin refining rule. The parameter ${\tau}_{s}$ is chosen as in (17). Accordingly, we remove arm $k^{*}$ from $A_{s}^{j}$ if there exists $k$ such that
$r_{k,s}^{t,j},r_{k^{*},s}^{t,j}>0\;\;\text{ and }\;\;\tilde{f}_{k}^{t}(x)-2r_{k,s}^{t,j}>\tilde{f}_{k^{*}}^{t}(x)+2r_{k^{*},s}^{t,j},\;\;\;x\in B_{s}^{j}\in\mathcal{B}_{t}.$ (29)
@@
Similar to (16), the first condition in (29) aims to ensure sufficient samples have been collected, since we notice $r_{k,s}^{t,j}>0$ implies at least one dataset provides $\log^{2}n$ samples.
@@
Input: Budgets $\varepsilon,\varepsilon_{1},\ldots,\varepsilon_{M}$, auxiliary sample sizes $n_{\mathrm{Q}_{1}},\ldots,n_{\mathrm{Q}_{m}}$.
Initialization: $\pi_{0}=\operatorname{Unif}([K])$, $\mathcal{B}_{0}=[0,1]^{d}$, $A_{0}^{0}=[K]$, $t=0$.
# jump-start
@@
                   Remove $k$ from $A_{s}^{j}$ if (29). # eliminating arms
                   if  $r_{k,s}^{t,j}<{\tau}_{s}$ for any $k$  then
                         $\mathcal{B}_{t}=\mathcal{B}_{t}\cup\{B_{s+1}^{2j-1},B_{s+1}^{2j}\}/B_{s}^{j}$. # refining bins
@@
            $t\leftarrow t+1$. $\mathcal{B}_{t}=\mathcal{B}_{t-1}$. Update $\pi_{t}$ by (8).
@@
# main loop
while True do
@@
             Remove $k$ from $A_{s}^{j}$ if (29). # eliminating arms
             if  $r_{k,s}^{t,j}<{\tau}_{s}$ for any $k$  then
                   $\mathcal{B}_{t}=\mathcal{B}_{t}\cup\{B_{s+1}^{2j-1},B_{s+1}^{2j}\}/B_{s}^{j}$. # refining bins
@@
      $t\leftarrow t+1$. $\mathcal{B}_{t}=\mathcal{B}_{t-1}$. Update $\pi_{t}$ by (8). Send $\pi_{t}$ to user $t$.
end while
Algorithm 3 LDP MAB with Auxiliary Data
@@ 4 Numerical experiments
We conduct experiments on both synthetic data (Section 4.1) and real-world data (Section 4.2), to respectively validate our theoretical findings and show superior performance of the proposed method. All experiments are conducted on a machine with 72-core Intel Xeon 2.60GHz and 128GB memory. Reproducible codes are available on GitHub [^1] https://github.com/Karlmyh/LDP-Contextual-MAB.
@@ 4.1 Simulation Studies
Synthetic Distribution
For distribution $\mathrm{P}$, we choose the marginal distribution $\mathrm{P}_{X}$ to be the uniform distribution on $\mathcal{X}=[0,1]^{d}$. For the reward function, let
@@
For better illustration, the reward functions are plotted in Figure 4. The auxiliary data distribution is taken as $\mathrm{Q}_{m,X}(x)=c_{norm}\|x-\mathrm{I}_{d}/2\|_{\infty}^{\gamma}$, where $\mathrm{I}_{d}$ is the $d$ dimensional vector with all entries equal to $1$. We can explicitly compute the normalizing constant $c_{norm}=2^{-\gamma}d/(d+\gamma)$. Figure 5 illustrates $\gamma=0.2,1,2$. The behavior policies for the source data are a discrete distribution with probability vector $\kappa/K+(2-2\kappa)/\{K(K-1)\}\cdot(0,\ldots,K-1)$ over $[K]$, which belongs to $\Lambda(K,\beta,\gamma,\kappa)$. In the experiments, we take $\varepsilon,\varepsilon_{m}\in\{1,2,4,8,1024\}$, covering commonly seen magnitudes of privacy budgets from low to high privacy regimes (Erlingsson et al., 2014; Apple, 2017) as well as the (essentially) non-private case.
@@
Figure 5: Illustration of marginal of source distributions.
An Illustration
We first illustrate how the auxiliary datasets benefit the learning process using a simple example. For $n_{\mathrm{P}}$ target samples, we consider the following metrics for $t\in[n_{\mathrm{P}}]$. For global performance, we use the overall averaged regret
@@
For local performance, we use two metrics at a fixed point $x$, the local averaged regret and the ratio of chosen arms:
@@
For the uniformly random policy, all three quantities should remain approximately unchanged for all time steps. For any effective policy, we expect to see $\overline{R}_{t}^{\mathrm{global}}(\pi)$ and $\overline{R}_{t}^{\mathrm{local}}(\pi,x)$ decreasing and $\overline{R}_{t}^{\mathrm{ratio}}(\pi,x,\pi^{*}(x))$ increasing. We use these average metrics instead of the additive regret because the zero-order trend is more apparent than the first-order trend for visualization. We consider three settings: learning without auxiliary data, with effective auxiliary data, and with weak auxiliary data. The results in Figure 6 show that auxiliary data significantly accelerates the learning process by eliminating sub-optimal arms in the early stages, effectively providing a jump-start that leads to faster descent in both local and global regret. Additionally, the quality of the auxiliary data determines the magnitude of this jump-start effect.
@@
Figure 6: We use $\varepsilon=1$, $n_{\mathrm{P}}=1000$, and $M=1$. The effective auxiliary data has $n_{\mathrm{Q}_{1}}=500$, $\varepsilon_{1}=8$, and $\gamma_{1}=0$. The weak auxiliary data has $n_{\mathrm{Q}_{1}}=500$, $\varepsilon_{1}=0.5$, and $\gamma_{1}=5$. Both auxiliary dataset has $\kappa_{1}=0$. We run a single trial as a showcase. The top row exhibits the global average regret curves. The middle row exhibits the local average regret curve at $x=(1/3,1/3)$. The bottom row exhibits the ratio of pulled arms at $x=(1/3,1/3)$, which is represented by the width of each color at the cross-section at the time $t$. Blue, orange, and green represent the arm 1, 2, and 3, respectively. Note that we know the best arm for $(1/3,1/3)$ is $1$, i.e., we expect to see the blue area increase. The black vertical lines indicate when one of the sub-optimal arms at $(1/3,1/3)$ is eliminated, leading to a phase transition in the local regret curves and arm ratios. It is observed that both types of auxiliary data bring forward the elimination, but the effective auxiliary data is significantly more impactful.
@@
(b) Regret curve with auxiliary data.
Figure 7: Regret curves with $\varepsilon\in\{1,2,4,8,1024\}$ and $n_{\mathrm{P}}$ between $10000$ and $160000$. In (b), we use auxiliary data with $n_{\mathrm{Q}_{1}}=5000$ and $\varepsilon_{1}=8$. The colored areas are 95% confidence intervals.
Sample Sizes
We first analyze the regret curve with respect to sample sizes in Figure 7. The regret increases in a sub-linear manner with respect to $n_{\mathrm{P}}$, while the growth trend becomes slower as $\varepsilon$ increases. This aligns with the theoretical finding in Theorem 2.6. Moreover, under the same $\varepsilon$, the growth trend is less steep with the participation of auxiliary data in Figure 7(b). Interestingly, we note that with auxiliary data, the confidence interval of non-private data ($\varepsilon=1024$) becomes wider since the high variance brought by the (privatized) auxiliary data becomes significant in this case. Similar observations are also observed in Figure 8, where we fix the target data to observe the improvements brought by auxiliary data under different settings. The improvements are more significant for smaller $\gamma$, larger $n_{\mathrm{Q}_{m}}$ and $\varepsilon_{m}$, i.e. when the auxiliary data has higher quality. This phenomenon is well explained by the regret rate characterized in Theorem 3.4. We also note that confidence intervals are much wider for small $\varepsilon$ and $\varepsilon_{m}$ in both Figure 7 and 8, due to the instability brought by the injected Laplacian noise.
(a) $\gamma=0,\varepsilon_{m}=1$
(b) $\gamma=0.2,\varepsilon_{m}=1$
(c) $\gamma=2,\varepsilon_{m}=1$
(d) $\gamma=0,\varepsilon_{m}=2$
(e) $\gamma=0.2,\varepsilon_{m}=2$
(f) $\gamma=2,\varepsilon_{m}=2$
Figure 8: Regret curve over $n_{\mathrm{Q}_{m}}$ with different $\gamma$ and $\varepsilon_{m}$, using $\varepsilon=1$, $n_{\mathrm{P}}=80000$, and $M=2$. We set $\kappa_{m}=0,m=1,2$. The colored areas are 95% confidence intervals.
Underlying Parameters
We proceed to investigate the roles of the underlying parameters that control the auxiliary data, namely $\kappa$ and $\gamma$. In the bottom of Figure 9(a), we observe that with large $\varepsilon_{m}$, the regret is notably decreasing with respect to $\kappa$. This aligns with the rate in (22). In contrast, when $\varepsilon_{m}$ is small, e.g. in the top panel of Figure 9(a), regret barely varies as $\kappa$ changes. This is explained by the observation that (22) is dominated by $\mathrm{P}$ data if $\varepsilon_{m}$ is too small. In this case, the auxiliary dataset does not affect the learning process much, and the variation due to $\kappa$ is negligible. For $\gamma$ in Figure 9(b), we observe a similar phenomenon, where the regret is increasing with respect to $\gamma$, while the slope is controlled by $\varepsilon_{m}$.
@@
Figure 9: (a) contains regret curve over $\kappa$ with different auxiliary budgets (above $\varepsilon_{m}=1$, below $\varepsilon_{m}=8$), using $M=2$, $\gamma=0.2$, $n_{\mathrm{P}}=40000$ and $\varepsilon=2$; (b) contains regret curve over $\gamma$ with different auxiliary budgets (above $\varepsilon_{m}=2$, below $\varepsilon_{m}=8$), using $M=2$, $\kappa_{m}=1$, $n_{\mathrm{P}}=40000$ and $\varepsilon=2$; (c) contains comparison between increasing and decreasing auxiliary privacy budgets with different sample size (above $n_{\mathrm{P}}=10000$, below $n_{\mathrm{P}}=80000$), using $M=2$ and $\varepsilon_{m}\in\{2,8\}$. The colored areas are 95% confidence intervals.
Order of Auxiliary Data
We demonstrate potential improvements by carefully arranging the order in which auxiliary datasets are introduced during the jump-start stage. We conduct two sets of experiments with $M=2$, where one auxiliary dataset has a small $\varepsilon_{m}$ (high-quality data), and the other has a large $\varepsilon_{m}$ (low-quality data). The sole difference between the two experiments lies in the order of the auxiliary datasets. In Figure 9(c) (top), a significant performance gap is observed between starting with high-quality data versus low-quality data. This gap arises due to arms that were mistakenly removed by low-quality data. Sometimes, the algorithm can be overly aggressive in eliminating arms. It may accidentally remove the true arm, leading to persistent regret in that area, even when the forthcoming target data is quite effective. In Figure 9(c) (bottom), the auxiliary data may even become detrimental ($\varepsilon=8$) when low-quality data is used first, whereas it remains beneficial when high-quality data is prioritized. These results suggest that starting with high-quality data is recommended for achieving better overall performance.

### v2

@@ 3.3 Upper Bound Methodology
$r_{k,s}^{t,j}:=\sqrt{\frac{C_{n}\sum_{m=0}^{M}(\lambda_{t,k,s}^{m,j})^{2}\left\{\left(\varepsilon_{m}^{-2}t_{s}^{m,j}\right)\vee\sum_{i=1}^{T_{m}(t)}\tilde{U}_{i,k,s}^{\mathrm{Q}_{m},j}a_{i,s}^{m,j}\right\}}{\left(\sum_{m=0}^{M}\lambda_{t,k,s}^{m,j}\sum_{i=1}^{T_{m}(t)}\tilde{U}_{i,k,s}^{\mathrm{Q}_{m},j}a_{i,s}^{m,j}\right)^{2}}},$ (28)
@@
where $C_{n}\asymp\log(n)$ with its exact expression specified in the proof. As shown in LABEL:lem:finitedifferenceboundauxilary, (28) provides a valid high-probability confidence bound for the multiple-source estimator, with a rationale similar to that of (14). Note that the term $\sum_{i=1}^{T_{m}(t)}\tilde{U}_{i,k,s}^{\mathrm{Q}_{m},j}a_{i,s}^{m,j}$ approximately corresponds to the number of samples in the $m$-th dataset falling in $B_{s}^{j}$ while pulling arm $k$. This quantity generally increases with $\kappa_{m}$ and decreases with $\gamma_{m}$, in view of the definitions of these quantities. Therefore, as a statistic, (28) naturally encodes information about $\kappa_{m}$ and $\gamma_{m}$, which is the key reason that enables our estimator and thus algorithm to be adaptive to these unknown parameters.
Given the newly designed local estimator (26) and the confidence bound (28), the algorithm can then conduct arm elimination and bin refining. In particular, an arm $k^{*}$ is removed from the active arm set $A_{s}^{j}$ of a bin $B_{s}^{j}\in\mathcal{B}_{t}$ if there exists $k\neq k^{*}$ such that
$r_{k,s}^{t,j},r_{k^{*},s}^{t,j}>0\;\;\text{ and }\;\;\tilde{f}_{k}^{t}(x)-2r_{k,s}^{t,j}>\tilde{f}_{k^{*}}^{t}(x)+2r_{k^{*},s}^{t,j}.$ (29)
@@
Similar to (16), the first condition in (29) aims to ensure that sufficient samples have been collected, since we notice $r_{k,s}^{t,j}>0$ implies at least one dataset provides $\log^{2}(n)$ samples. A bin $B_{s}^{j}\in\mathcal{B}_{t}$ is refined if $r_{k,s}^{t,j}<{\tau}_{s}$ for some $k\in A_{s}^{j}$, where the parameter ${\tau}_{s}$ is set as in (17).
@@
Input:Budgets $\varepsilon,\varepsilon_{1},\ldots,\varepsilon_{M}$, auxiliary sample sizes $n_{\mathrm{Q}_{1}},\ldots,n_{\mathrm{Q}_{m}}$, target sample size $n_{\mathrm{P}}$.
Initialization: $\pi_{1}=\operatorname{Unif}([K])$, $\mathcal{B}_{1}=\{B_{0}^{1}\}=\{[0,1]^{d}\}$, $A_{0}^{1}=[K]$, $t=1$.
# jump-start via auxiliary data
@@
                   Remove $k$ from $A_{s}^{j}$ if (29) holds. # eliminating arms
                   if  $r_{k,s}^{t,j}<{\tau}_{s}$ for some $k\in A_{s}^{j}$  then
                         $\mathcal{B}_{t}=\mathcal{B}_{t}\cup\{B_{s+1}^{2j-1},B_{s+1}^{2j}\}\setminus B_{s}^{j}$. # refining bins
@@
            Set $t\leftarrow t+1$, $\mathcal{B}_{t}=\mathcal{B}_{t-1}$ and update $\pi_{t}$ by (8).
@@
# interaction on target data
for $i\in[n_{\mathrm{P}}]$ do
       The user $i$ receives $\pi_{t}$ from the server, pulls an arm via $\pi_{t}$ and receives the reward.
@@
             Remove $k$ from $A_{s}^{j}$ if (29) holds. # eliminating arms
             if  $r_{k,s}^{t,j}<{\tau}_{s}$ for some $k\in A_{s}^{j}$  then
                   $\mathcal{B}_{t}=\mathcal{B}_{t}\cup\{B_{s+1}^{2j-1},B_{s+1}^{2j}\}\setminus B_{s}^{j}$. # refining bins
@@
      Set $t\leftarrow t+1$ and $\mathcal{B}_{t}=\mathcal{B}_{t-1}$. Update $\pi_{t}$ by (8) and send to the next user.
end for
(For simplicity, we do not explicitly separate the user and server sides in the presentation.)
Algorithm 3 The nonparametric MAB algorithm under LDP with auxiliary data
(For simplicity, we do not explicitly separate the user and server sides in the presentation.)
@@ 4 Numerical experiments
In this section, we conduct numerical experiments on both synthetic data (Section 4.1) and real-world data (Section 4.2), to respectively validate our theoretical findings and show promising performance of the proposed method. All experiments are conducted on a machine with 72-core Intel Xeon 2.60GHz and 128GB memory. Reproducible codes are available on GitHub [^1] https://github.com/Karlmyh/LDP-Contextual-MAB.
@@ 4.1 Simulation Studies
Synthetic Distributions. For distribution $\mathrm{P}$, we choose the marginal distribution $\mathrm{P}_{X}$ to be the uniform distribution on $\mathcal{X}=[0,1]^{d}$. For the reward function, let
@@
The reward functions are plotted in Figure 4. The auxiliary data distribution is taken as $\mathrm{Q}_{m,X}(x)=c_{norm}\|x-\mathrm{I}_{d}/2\|_{\infty}^{\gamma}$, where $\mathrm{I}_{d}$ is the $d$ dimensional vector with all entries equal to $1$. We can explicitly compute the normalizing constant $c_{norm}=2^{-\gamma}d/(d+\gamma)$. Figure 5 illustrates $\gamma=0.2,1,2$. The behavior policies for the auxiliary data are a discrete distribution with probability vector $\kappa/K+(2-2\kappa)/\{K(K-1)\}\cdot(0,\ldots,K-1)$ over $[K]$, which belongs to $\Lambda(K,\beta,\gamma,\kappa)$.
In the numerical experiments, we fix $K=3$ and take $\varepsilon,\varepsilon_{m}\in\{1,2,4,8,1024\}$, covering commonly seen magnitudes of privacy budgets from high to low privacy regimes (Erlingsson et al.,, 2014; Apple,, 2017) as well as the (essentially) non-private case. To conserve space, the implementation details of all algorithms can be found in LABEL:sec:implementationdetail of the supplement. In LABEL:sec:add_exp of the supplement, we further provide numerical results under an alternative simulation setting with more complex reward functions, where similar findings as the ones seen below in Figures 7-9 are observed. All simulation results presented below are based on 100 repetitions unless otherwise noted.
@@
Figure 5: Illustration of marginal distribution $Q_{m,X}$ of source data.
An Illustrative Example. We first illustrate how the auxiliary datasets benefit the learning process via a simple example. For $n_{\mathrm{P}}$ target samples, we consider the following metrics for $t\in[n_{\mathrm{P}}]$. For global performance, we use the overall averaged regret
@@
For local performance, we use two metrics at a fixed point $x\in\mathcal{X}$, the local averaged regret and the ratio of chosen arms:
@@
For a naive policy that selects arms uniformly at random, all three quantities should remain approximately unchanged for all time steps. For any effective policy, we expect to see $\overline{R}_{t}^{\mathrm{global}}(\pi)$ and $\overline{R}_{t}^{\mathrm{local}}(\pi,x)$ decreasing and $\overline{R}_{t}^{\mathrm{ratio}}(\pi,x,\pi^{*}(x))$ increasing over time. We use the average metrics instead of cumulative regret as the zero-order trend is more apparent than the first-order trend for visualization. We consider three settings: learning without auxiliary data, with effective auxiliary data, and with weak auxiliary data. The results in Figure 6 show that auxiliary data significantly accelerates the learning process by eliminating sub-optimal arms in the early stages, effectively providing a jump-start that leads to faster descent in both local and global regret. Additionally, the quality of the auxiliary data determines the magnitude of this jump-start effect.
@@
Figure 6: We set $\varepsilon=1$, $n_{\mathrm{P}}=1000$, and $M=1$. The effective auxiliary data has $n_{\mathrm{Q}_{1}}=500$, $\varepsilon_{1}=8$, and $\gamma_{1}=0$. The weak auxiliary data has $n_{\mathrm{Q}_{1}}=500$, $\varepsilon_{1}=0.5$, and $\gamma_{1}=5$. Both auxiliary dataset has $\kappa_{1}=1$. We run a single trial as a showcase. The top row exhibits the global average regret curves. The middle row exhibits the local average regret curve at $x=(1/3,1/3)$. The bottom row exhibits the ratio of pulled arms at $x=(1/3,1/3)$, which is represented by the width of each color at the cross-section at the time $t$. Blue, orange, and green represent the arm 1, 2, and 3, respectively. Note that we know the best arm for $(1/3,1/3)$ is $1$, i.e., we expect to see the blue area increase. The black vertical lines indicate when one of the sub-optimal arms at $(1/3,1/3)$ is eliminated, leading to a phase transition in the local regret curves and arm ratios. It is observed that both types of auxiliary data bring forward the elimination of sub-optimal arms (such an event is marked by vertical dashed line), but the effective auxiliary data is significantly more impactful.
@@
(b) Regret curve with auxiliary data
Figure 7: Regret with $\varepsilon\in\{1,2,4,8,1024\}$ and $n_{\mathrm{P}}\in\{1,2,4,6,8,12,16\}\times 10^{4}$. In (b), we use auxiliary data with $n_{\mathrm{Q}_{1}}=5000$, $\varepsilon_{1}=8$, $\gamma_{1}=0$ and $\kappa_{1}=1$. The colored areas are 95% confidence intervals.
Sample Sizes. We first analyze the regret curve with respect to sample sizes $n_{\mathrm{P}}$ in Figure 7. The regret increases in a sub-linear manner with respect to $n_{\mathrm{P}}$, while the growth trend becomes slower as $\varepsilon$ increases. This aligns with the theoretical finding in Theorem 2.6. Moreover, under the same $\varepsilon$, the growth trend is less steep with the participation of auxiliary data in Figure 7(b). Interestingly, we note that with auxiliary data, the confidence interval of non-private data ($\varepsilon=1024$) becomes wider since the high variance brought by the (privatized) auxiliary data becomes significant in this case. A similar phenomenon is also observed in Figure 8, where we fix the sample size of the target data to examine the improvements brought by auxiliary data under different settings. As expected, the improvements are more notable for smaller $\gamma$, larger $n_{\mathrm{Q}_{m}}$ and $\varepsilon_{m}$, i.e. when the auxiliary data has higher quality. This phenomenon is well explained by the regret rate characterized in Theorem 3.4. We also note that confidence intervals are much wider for small $\varepsilon$ and $\varepsilon_{m}$ in both Figures 7 and 8, due to the high variance of the injected Laplacian noise.
(a) $\gamma=0,\varepsilon=1$
(b) $\gamma=0.2,\varepsilon=1$
(c) $\gamma=2,\varepsilon=1$
(d) $\gamma=0,\varepsilon=2$
(e) $\gamma=0.2,\varepsilon=2$
(f) $\gamma=2,\varepsilon=2$
Figure 8: Regret curves over $n_{\mathrm{Q}_{m}}\in\{0,1,2,5,10\}\times 10^{4}$ at different $(\gamma,\varepsilon_{m})$, while we fix $n_{\mathrm{P}}=80000$, and fix $M=2$ and $\kappa_{1}=\kappa_{2}=1$. The colored areas are 95% confidence intervals.
Underlying Parameters. We proceed to investigate the roles of the underlying parameters that control the quality of the auxiliary data, namely $\kappa$ and $\gamma$. In the bottom panel of Figure 9(a), we observe that with large $\varepsilon_{m}$, the regret is notably decreasing with respect to $\kappa$. This aligns with the regret upper bound in (22). In contrast, when $\varepsilon_{m}$ is small, e.g. in the top panel of Figure 9(a), regret barely varies as $\kappa$ changes. This is explained by the observation that (22) is dominated by the target data if $\varepsilon_{m}$ is too small. In this case, the auxiliary dataset does not affect the learning process much, and the variation due to $\kappa$ is negligible. For $\gamma$ in Figure 9(b), we observe a similar phenomenon, where the regret is increasing with respect to $\gamma$, while the slope is controlled by $\varepsilon_{m}$.
@@
Figure 9: (a) Regret curves over $\kappa\in\{0,0.2,\cdots,0.8,1\}$ with different auxiliary privacy budgets (top $\varepsilon_{m}=1$, bottom $\varepsilon_{m}=8$), while fixing $M=2$, $\gamma=0.2$, $n_{\mathrm{P}}=40000$ and $\varepsilon=2$; (b) Regret curves over $\gamma\in\{0,0.2,\cdots,1.8,2\}$ with different auxiliary privacy budgets (top $\varepsilon_{m}=2$, bottom $\varepsilon_{m}=8$), while fixing $M=2$, $\kappa=1$, $n_{\mathrm{P}}=40000$ and $\varepsilon=2$; (c) Comparison of regret curves when the two auxiliary datasets enter the jump-start stage in different orders, for different target data budgets $\varepsilon\in\{1,2\}$ (top $n_{\mathrm{P}}=10000$, bottom $n_{\mathrm{P}}=80000$). The colored areas are 95% confidence intervals.
Order of Auxiliary Data. We demonstrate potential improvements by carefully arranging the order in which auxiliary datasets are introduced during the jump-start stage. We conduct two sets of experiments with $M=2$, where one auxiliary dataset has a small $\varepsilon_{m}=2$ (low-quality data), and the other has a large $\varepsilon_{m}=8$ (high-quality data). The only difference between the two experiments lies in which of the two auxiliary datasets enters the jump-start stage first. In Figure 9(c), a significant performance gap on the target data is observed between starting with high-quality auxiliary data versus starting with low-quality data. We believe this gap arises due to arms that were mistakenly removed by low-quality auxiliary data. In particular, the algorithm can sometimes be overly aggressive in eliminating arms during the jump-start stage, which may incorrectly remove the optimal arm, leading to persistent regret in that area for the target data. These results suggest that starting with high-quality auxiliary data is recommended for achieving better overall performance.